# On degenerate metric resolvent: well-definedness and new convergence analysis

### v3

@@ 3.3 Examples
We then deduce that $\mathsf{ran}(\mathcal{A}+\mathcal{Q})=[-1,0]\times\mathbb{R}$. Clearly, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\supseteq\mathsf{ran}\mathcal{Q}=\{0\}\times\mathbb{R}$. By Theorem 3.1, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ must be maximally monotone. Indeed, if the input $y=0$, we then have $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}:0\mapsto\{0\}\times\mathbb{R}$. Then, $\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})\supseteq(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})(0)=(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})(0)=\{0\}\times\mathbb{R}=\mathsf{ran}\mathcal{Q}$, and thus by restricted Minty’s theorem, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
However, $\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A}=[0,1]\times\mathbb{R}$, $0\notin\mathsf{sri}(\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q})$. This shows that the condition of $0\in\mathsf{sri}(\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q})$ is sufficient but not necessary condition for the full domain of $\mathcal{T}$.
@@ 4.1 Restricted injectivity
As mentioned in Sect. 1.2, a standard condition for the single-valuedness of $\mathcal{T}$ is disjoint injectivity of $\mathcal{A}+\mathcal{Q}$ [5, Theorem 2.1-(ix)], [13, Proposition 3.8]. Moreover, a sufficient and necessary condition, proposed in [12, Proposition 2.1-(2)], is that $\mathcal{A}+\mathcal{Q}$ is injective on $\mathsf{ran}\mathcal{Q}$. Here, Theorem 4.2 claims that this condition can be further weakened to to the disjoint injectivity on $\mathsf{ran}\mathcal{Q}$ w.r.t. $\mathsf{ker}\mathcal{Q}$ only, under Assumptions 1 and 2.
@@
We say that $\mathcal{B}$ is disjointly injective on $\mathsf{ran}\mathcal{Q}$ w.r.t. $\mathsf{ker}\mathcal{Q}$, if
$\mathcal{B}y_{1}\cap\mathcal{B}y_{2}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing\Longrightarrow y_{1,\mathsf{k}}=y_{2,\mathsf{k}},\quad\forall y_{1},y_{2}\in\mathsf{dom}\mathcal{B}.$ (8)
@@
Under Assumptions 1 and 2, for any given input $x_{\mathsf{r}}\in\mathsf{ran}\mathcal{Q}\cap\mathsf{dom}\mathcal{T}$, $y\in\mathcal{T}x_{\mathsf{r}}$ is single-valued, if and only if one of the following equivalent conditions holds:
(i) $\mathcal{A}+\mathcal{Q}$ is disjointly injective on $\mathsf{ran}\mathcal{Q}$ w.r.t. $\mathsf{ker}\mathcal{Q}$;
(ii) For any input $x_{\mathsf{r}}\in\mathsf{ran}\mathcal{Q}$, given the corresponding $y_{\mathsf{r}}=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})$, the constrained inclusion problem of $\mathcal{A}(y_{\mathsf{r}}+y_{\mathsf{k}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$, s.t. $\mathcal{Q}y_{\mathsf{k}}=0$, has a unique solution of $y_{\mathsf{k}}$;
(iii) Suppose there exists a linear, bounded and surjective operator $\mathcal{E}:\mathcal{H}\mapsto D^{\prime}$, for some real Hilbert space $D^{\prime}$, such that $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}=\mathcal{E}^{\top}\mathcal{E}$. Then, for any input $x_{\mathsf{r}}\in\mathsf{ran}\mathcal{Q}$ given the corresponding $y_{\mathsf{r}}=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})$, the inclusion problem $\mathcal{A}(y_{\mathsf{r}}+\mathcal{E}^{\top}b)\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$ has a unique solution of $b\in D^{\prime}$;
(iv) $\mathsf{ker}\mathcal{A}\cap(\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})-y)=\{0\}$.
@@
Single-$\Longrightarrow$(i): Let us prove (8) by contradiction. If (8) does not hold, it implies that $\exists y_{1},y_{2}\mathsf{dom}\mathcal{A}$, such that $(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$, and $y_{1,\mathsf{k}}\neq y_{2,\mathsf{k}}$. Then, take $x$, such that $\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}$, i.e., $\{y_{1},y_{2}\}\subseteq\mathcal{T}x$. It has been shown in Lemma 2.3 that $y_{1,\mathsf{r}}=y_{2,\mathsf{r}}$. Combining with $y_{1,\mathsf{k}}\neq y_{2,\mathsf{k}}$, we obtain $y_{1}\neq y_{2}$. This means that $\mathcal{T}$ is not single-valued, which is contradiction.
@@
$y\in\mathcal{T}x\Longleftrightarrow\mathcal{A}(y_{\mathsf{r}}+y_{\mathsf{k}})=\mathcal{A}y\owns\mathcal{Q}(x-y)=\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}}),$ (9)
which results in
$y_{\mathsf{r}}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}}),\quad y_{\mathsf{k}}\in\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}}).$ (10)
@@
(i)$\Longleftrightarrow$(ii): Since $(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$, pick a point $x$, such that
@@
Under Assumptions 1 and 2, $\mathcal{T}$ is single-valued, if $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is single-valued.
@@
(9) can be rewritten as $y_{\mathsf{r}}+y_{\mathsf{k}}\in\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$. Then, apply the projection $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}$ to both sides yields $y_{\mathsf{k}}\in\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$, which completes the proof.
@@
(i) The condition of Corollary 4.4 is sufficient, but not necessary. Actually, the single-valuedness of $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is much stronger than Theorem 4.2-(iv): Even if $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is multi-valued, $y_{\mathsf{k}}$ could still be uniquely determined by (9). This depends on the interaction between $\mathcal{A}$ and $\mathcal{Q}$.
(ii) There is no need to enforce $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ to be single-valued, since the uniqueness of $y_{\mathsf{r}}$ has been guaranteed by Lemma 2.3. This is irrelevant to single-valuedness of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$, which is, indeed, often multi-valued in practice.
(iii) A big merit of condition Theorem 4.2-(iii), compared to (ii), is that the unconstrained version is easier to use and solve.
(iv) Theorem 4.2-(iv) has similar spirit with the most recent result of [17, Theorem 3.1], for regularized optimization problems. Their potential connections deserve further explorations in future work.
We present the following result as summary for the well-definedness of $\mathcal{T}$.
@@ 4.2 More remarks and examples
The following guiding examples show how the single-valuedness of $\mathcal{T}$ depends on the choice of $\mathcal{Q}$. Given a function $f:\mathbb{R}^{2}\mapsto\mathbb{R}:(x_{1},x_{2})\mapsto|x_{1}|$, consider a minimization problem w.r.t. $x=(x_{1},x_{2})\in\mathbb{R}^{2}$:
@@
$\begin{bmatrix}0\\ 0\end{bmatrix}\in\begin{bmatrix}\partial|\cdot|&0\\ 0&0\end{bmatrix}\begin{bmatrix}x_{1}^{\star}\\ x_{2}^{\star}\end{bmatrix}+\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\begin{bmatrix}x_{1}^{\star}-b\\ x_{2}^{\star}-0\end{bmatrix}$
with the input variable $y=(b,0)$.
Obviously, the optimal solution set is $X^{\star}=\{x_{1}^{\star}\}\times\mathbb{R}$, where $x_{1}^{\star}$ is given by a well-known soft-thresholding function $S(\cdot)$ of $b$:
$x_{1}^{\star}=S(b):=\left\{\begin{array}[]{ll}b-1,&\textrm{if\ }b>1;\\ b+1,&\textrm{if\ }b<-1;\\ 0,&\textrm{if\ }|b|\leq 1.\end{array}\right.$
One can easily check that neither Theorem 4.2 nor Corollary 4.4 is satisfied. For instance,
$\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(y_{\mathsf{r}}-x^{\star}_{\mathsf{r}})=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(y_{\mathsf{r}}-x^{\star}_{\mathsf{r}})=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\begin{bmatrix}b-S(b)\\ 0\end{bmatrix}=\{0\}\times\mathbb{R},$
which is multi-valued.
@@
$\begin{bmatrix}0\\ 0\end{bmatrix}\in\begin{bmatrix}\partial|\cdot|&0\\ 0&0\end{bmatrix}\begin{bmatrix}x_{1}^{\star}\\ x_{2}^{\star}\end{bmatrix}+\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\begin{bmatrix}x_{1}^{\star}-0\\ x_{2}^{\star}-b\end{bmatrix}$
with input $y=(0,b)$. The unique solution is obviously $x^{\star}=(0,b)$. On the other hand, it is easy to check that $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(y_{\mathsf{r}}-x^{\star}_{\mathsf{r}})=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\begin{bmatrix}0\\ 0\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}$. This shows, by Corollary 4.4, the uniqueness of solution, $\forall b\in\mathbb{R}$.

### v4

@@ 3.3 Examples
We then deduce that $\mathrm{ran}(\mathcal{A}+\mathcal{Q})=[-1,0]\times\mathbb{R}$. Clearly, $\mathrm{ran}(\mathcal{A}+\mathcal{Q})\supseteq\mathrm{ran}\mathcal{Q}=\{0\}\times\mathbb{R}$. By Theorem 3.1, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ must be maximally monotone. Indeed, if the input $y=0$, we then have $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}:0\mapsto\{0\}\times\mathbb{R}$. Then, $\mathrm{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}})\supseteq(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}})(0)=(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}})(0)=\{0\}\times\mathbb{R}=\mathrm{ran}\mathcal{Q}$, and thus by restricted Minty’s theorem, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ is maximally monotone.
However, $\mathrm{ran}\mathcal{Q}-\mathrm{ran}\mathcal{A}=[0,1]\times\mathbb{R}$, $0\notin\mathrm{sri}(\mathrm{ran}\mathcal{A}-\mathrm{ran}\mathcal{Q})$. This shows that the condition of $0\in\mathrm{sri}(\mathrm{ran}\mathcal{A}-\mathrm{ran}\mathcal{Q})$ is sufficient but not necessary condition for the full domain of $\mathcal{T}$.
@@ 4.1 Restricted injectivity
As mentioned in Sect. 1.2, a standard condition for the single-valuedness of $\mathcal{T}$ is disjoint injectivity of $\mathcal{A}+\mathcal{Q}$ [6, Theorem 2.1-(ix)], [18, Proposition 3.8]. Moreover, a sufficient and necessary condition, proposed in [17, Proposition 2.1-(2)], is that $\mathcal{A}+\mathcal{Q}$ is injective on $\mathrm{ran}\mathcal{Q}$. Here, Theorem 4.2 claims that this condition can be further weakened to to the disjoint injectivity on $\mathrm{ran}\mathcal{Q}$ w.r.t. $\mathrm{ker}\mathcal{Q}$ only, under Assumptions 1 and 2.
@@
The operator $\mathcal{B}$ is said to be disjointly injective on $\mathrm{ran}\mathcal{Q}$ w.r.t. $\mathrm{ker}\mathcal{Q}$, if
$\mathcal{B}y_{1}\cap\mathcal{B}y_{2}\cap\mathrm{ran}\mathcal{Q}\neq\varnothing\Longrightarrow y_{1,\mathsf{k}}=y_{2,\mathsf{k}},\quad\forall y_{1},y_{2}\in\mathrm{dom}\mathcal{B}.$ (9)
@@
Under Assumptions 1 and 2, for any given input $x_{\mathsf{r}}\in\mathrm{ran}\mathcal{Q}\cap\mathrm{dom}\mathcal{T}$, $y\in\mathcal{T}x_{\mathsf{r}}$ is single-valued, if and only if one of the following equivalent conditions holds:
(i) $\mathcal{A}+\mathcal{Q}$ is disjointly injective on $\mathrm{ran}\mathcal{Q}$ w.r.t. $\mathrm{ker}\mathcal{Q}$;
(ii) For any input $x_{\mathsf{r}}\in\mathrm{ran}\mathcal{Q}$, given the corresponding $y_{\mathsf{r}}=(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})$, the constrained inclusion problem of $\mathcal{A}(y_{\mathsf{r}}+y_{\mathsf{k}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$, s.t. $\mathcal{Q}y_{\mathsf{k}}=0$, has a unique solution of $y_{\mathsf{k}}$;
(iii) Suppose there exists a linear, bounded and surjective operator $\mathcal{E}:\mathcal{H}\mapsto D^{\prime}$, for some real Hilbert space $D^{\prime}$, such that $\mathcal{P}_{\mathrm{ker}\mathcal{Q}}=\mathcal{E}^{\top}\mathcal{E}$. Then, for any input $x_{\mathsf{r}}\in\mathrm{ran}\mathcal{Q}$ given the corresponding $y_{\mathsf{r}}=(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})$, the inclusion problem $\mathcal{A}(y_{\mathsf{r}}+\mathcal{E}^{\top}b)\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$ has a unique solution of $b\in D^{\prime}$;
(iv) $\mathrm{ker}\mathcal{Q}\cap(\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})-y)=\{0\}$.
@@
Single-$\Longrightarrow$(i): Let us prove (9) by contradiction. If (9) does not hold, it implies that $\exists y_{1},y_{2}\in\mathrm{dom}\mathcal{A}$, such that $(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}\cap\mathrm{ran}\mathcal{Q}\neq\varnothing$, and $y_{1,\mathsf{k}}\neq y_{2,\mathsf{k}}$. Then, take $x$, such that $\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}$, i.e., $\{y_{1},y_{2}\}\subseteq\mathcal{T}x$. It has been shown in Lemma 2.4 that $y_{1,\mathsf{r}}=y_{2,\mathsf{r}}$. Combining with $y_{1,\mathsf{k}}\neq y_{2,\mathsf{k}}$, we obtain $y_{1}\neq y_{2}$. This means that $\mathcal{T}$ is not single-valued, which is contradiction.
@@
$y\in\mathcal{T}x\Longleftrightarrow\mathcal{A}(y_{\mathsf{r}}+y_{\mathsf{k}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})\Longleftrightarrow\left\{\begin{array}[]{l}y_{\mathsf{r}}\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}}),\\ y_{\mathsf{k}}\in\mathcal{P}_{\mathrm{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}}).\end{array}\right.$ (10)
@@
(i)$\Longleftrightarrow$(ii): Since $(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}\cap\mathrm{ran}\mathcal{Q}\neq\varnothing$, pick a point $x$, such that
@@
Under Assumptions 1 and 2, $\mathcal{T}$ is single-valued, if $\mathcal{P}_{\mathrm{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is single-valued.
@@
(10) can be rewritten as $y_{\mathsf{r}}+y_{\mathsf{k}}\in\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$. Then, apply the projection $\mathcal{P}_{\mathrm{ker}\mathcal{Q}}$ to both sides yields $y_{\mathsf{k}}\in\mathcal{P}_{\mathrm{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$, which completes the proof.
@@
(i) The condition of Corollary 4.4 is sufficient, but not necessary. Actually, the single-valuedness of $\mathcal{P}_{\mathrm{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is much stronger than Theorem 4.2-(iv): Even if $\mathcal{P}_{\mathrm{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is multi-valued, $y_{\mathsf{k}}$ could still be uniquely determined by (10). This depends on not only the interaction between $\mathcal{A}$ and $\mathcal{Q}$, but the specific position of $x_{\mathsf{r}}$ in terms of geometry as well.
(ii) There is no need to enforce $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ to be single-valued, since the uniqueness of $y_{\mathsf{r}}$ has been guaranteed by Lemma 2.4. The uniqueness of $y_{\mathsf{r}}$ is irrelevant to single-valuedness of $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$, which is, indeed, often multi-valued in practice.
(iii) A big merit of condition Theorem 4.2-(iii), compared to (ii), is that the unconstrained version is easier to solve.
(iv) Theorem 4.2-(iv) has similar spirit with the most recent result of [26, Theorem 3.1], for regularized optimization problems. Their potential connections, particularly in the geometric aspect, deserve further explorations in future work.
The well-definedness of $\mathcal{T}$ is summarized below.
@@ 4.2 More remarks and examples
The interaction between $\mathcal{A}$ and $\mathcal{Q}$ also influences the single-valuedness of $\mathcal{T}$. The following guiding examples show how the single-valuedness of $\mathcal{T}$ depends on the choice of $\mathcal{Q}$, given the same $\mathcal{A}$. Given a function $f:\mathbb{R}^{2}\mapsto\mathbb{R}:(x_{1},x_{2})\mapsto|x_{1}|$, consider a minimization problem w.r.t. $x=(x_{1},x_{2})\in\mathbb{R}^{2}$:
@@
$\begin{bmatrix}0\\ 0\end{bmatrix}\in\underbrace{\begin{bmatrix}\partial|\cdot|&0\\ 0&0\end{bmatrix}}_{\mathcal{A}}\begin{bmatrix}x_{1}^{\star}\\ x_{2}^{\star}\end{bmatrix}+\underbrace{\begin{bmatrix}1&0\\ 0&0\end{bmatrix}}_{\mathcal{Q}_{1}}\Bigg{(}\begin{bmatrix}x_{1}^{\star}\\ x_{2}^{\star}\end{bmatrix}-\underbrace{\begin{bmatrix}b\\ 0\end{bmatrix}}_{y}\Bigg{)}.$
It is easy to check that $0\in\mathrm{ri}(\mathrm{ran}\mathcal{A}-\mathrm{ran}\mathcal{Q}_{1})=\mathbb{R}\times\{0\}$, and by Corollary 3.7, (15) always has solutions for any $b\in\mathbb{R}$. Actually, the optimal solution set is $X^{\star}=\{x_{1}^{\star}\}\times\mathbb{R}$, where $x_{1}^{\star}$ is given by a well-known soft-thresholding function $S_{\lambda}(\cdot)$ of $b$, with the threshold $\lambda=1$ [53].
One can easily check that neither Theorem 4.2 nor Corollary 4.4 is satisfied. For instance, $\mathcal{P}_{\mathrm{ker}\mathcal{Q}_{1}}\mathcal{A}^{-1}\mathcal{Q}_{1}(y_{\mathsf{r}}-x^{\star}_{\mathsf{r}})=\mathcal{P}_{\mathrm{ker}\mathcal{Q}_{1}}\mathcal{A}^{-1}\begin{bmatrix}b-S_{1}(b)\\ 0\end{bmatrix}=\{0\}\times\mathbb{R}$, which is multi-valued.
@@
$\begin{bmatrix}0\\ 0\end{bmatrix}\in\underbrace{\begin{bmatrix}\partial|\cdot|&0\\ 0&0\end{bmatrix}}_{\mathcal{A}}\begin{bmatrix}x_{1}^{\star}\\ x_{2}^{\star}\end{bmatrix}+\underbrace{\begin{bmatrix}0&0\\ 0&1\end{bmatrix}}_{\mathcal{Q}_{2}}\Bigg{(}\begin{bmatrix}x_{1}^{\star}\\ x_{2}^{\star}\end{bmatrix}-\underbrace{\begin{bmatrix}0\\ b\end{bmatrix}}_{y}\Bigg{)}.$
It is easy to check $0\in\mathrm{ri}(\mathrm{ran}\mathcal{A}-\mathrm{ran}\mathcal{Q}_{2})=[-1,1]\times\mathbb{R}$, and $\mathcal{P}_{\mathrm{ker}\mathcal{Q}_{2}}\mathcal{A}^{-1}\mathcal{Q}_{2}(y_{\mathsf{r}}-x^{\star}_{\mathsf{r}})=\mathcal{P}_{\mathrm{ker}\mathcal{Q}_{2}}\mathcal{A}^{-1}\begin{bmatrix}0\\ 0\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}$. This demonstrates, by Corollary 4.4, the uniqueness of solution for any $b\in\mathbb{R}$. Indeed, the unique solution is obviously $x^{\star}=(0,b)$.